# A Calculus of Inheritance

### 10 Mar 2026 v6

@@ 6.1. Emergent Phenomena
In developing real programs in MIXINv2, we encountered three phenomena that we did not design for. They can be seen in the source code (Yang, 2026) and the Case Study (Section 5). Below we explain why they arise from the three primitives.
@@
Single-valued computation and set-valued computation are the same mechanism. The semantic equations of Section 3 resolve every property to a set of paths, so operations defined on single values automatically distribute over all combinations of set-valued inputs (Section 5). When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise. [^9] Inheritance-calculus to relational algebra: path from root to leaf == tuple; trie union == relation (set of tuples); $\mathtt{bases}$ (equation 4) == cross-join; $\mathtt{supers}$ (equation 2) == union; $\mathtt{properties}$ (equation 1) == projection. The carrier is a trie rather than a flat relation, and observation is coinductive: the observer drives computation by querying progressively deeper paths. This contrasts with Datalog’s bottom-up materialization, but the relational structure of the computation is the same. Appendix F makes this correspondence precise with a systematic encoding of Datalog into inheritance-calculus, demonstrating that variable joins, constant matches, multi-body rules, and recursive rules are all expressible using only the three primitives of Section 2.
@@ 6.2. Semantic Alternatives
The preceding item established that inheritance-calculus is immune to nonextensibility. We now ask whether an alternative design could share this property by surveying every candidate known to us.
@@
The analysis below assumes that the AST is a named tree:every edge from parent to child carries a label, so that every node is uniquely identified by the path from the root. For configuration languages this is naturally the case: every key in a JSON, YAML, or Nix attribute set is a label, and nesting produces paths. [^12] Nix is a functional language with first-class functions; its attribute sets alone do not form a named-tree DSL. The NixOS module system (NixOS Contributors, 2024), however, restricts the user-facing interface to nested attribute sets composed by recursive merging, and it is this DSL-level structure that naturally forms a named tree. For mainstream programming languages the AST is not directly a named tree, since anonymous expression positions and implicit scopes lack labels. After ANF transformation (Flanagan et al., 1993), however, every intermediate result is bound to a name, and the remaining anonymous positions [^13] For example, the body of a $\mathbf{let}$. can be assigned synthetic fresh labels such as $\mathrm{result}$ and $\mathrm{tailCall}$ in Section 4, yielding a named tree.
@@
Scala (multi-paradigm) class, trait, object, method, block, conditional, match, while, throw, … variable, member access, extends, with, implicits, import, …
@@
Only the parameters placed by the author are extension points; the function body is closed. To make a new aspect extensible, the original function must be rewritten to accept an additional parameter. This applies to every language whose composition mechanism is function application, whether functional, object-oriented, or multi-paradigm, because method calls, virtual dispatch, and type-class resolution all reduce to $\beta$-reduction. The Expression Problem (Wadler, 1998) is hard in all of them for the same reason: extensibility is opt-in. The frameworks that simulate immunity to nonextensibility (visitor pattern, object algebras (d. S. Oliveira and Cook, 2012), finally tagless interpreters (Carette et al., 2009)) are the machinery needed to work around this limitation.
@@
Removing or conditionally including definitions can introduce cycles in the dependency graph, causing the recursive evaluation (Appendix A) to diverge.
@@
Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (B.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution. However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason. When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route. The NixOS module system (NixOS Contributors, 2024) raises a duplicate-declaration error (Appendix D). This rejects nested extensions where two independently authored modules define the same option, a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
Our first three implementations of MIXINv2 used standard techniques for self-reference resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural. A closure captures a single environment, so a fixed-point combinator solves for a single self-reference. A stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-path situation that multiple inheritance naturally produces (Appendix D). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
@@
Single-target $\mathtt{this}$ nested extensions rejected NixOS modules (NixOS Contributors, 2024)
@@
The semantics of inheritance-calculus was not chosen from a menu. We arrived at it by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-path compositions that the Expression Problem demands. [^15] Three successive implementations confirmed this; see the discussion above. Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
@@ 7.1. Computational Models and $\lambda$-Calculus Semantics
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II (II is redundant: $I=SKK$). The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus. Semantically, the two share the same models: Meyer (Meyer, 1982) showed that combinatory algebras satisfying five additional conditions are exactly lambda algebras. No independent semantic domain for SKI has ever been exhibited from which $\lambda$-calculus semantics could be derived. The direction has always been the reverse: Scott’s $D_{\infty}$ was built for the $\lambda$-calculus, and SKI was interpreted inside it afterwards. Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine. The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements. Each of these models achieves Turing completeness by removing structure from an existing model: fewer abstractions, worse ergonomics, no new expressiveness. None of these reductions exhibits emergent phenomena absent from the model it reduces.
@@
The semantics of the $\lambda$-calculus can be defined in several fundamentally different ways. Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984). This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991). It also requires a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990). Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem. Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972). The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes (Plotkin, 1975). Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995) at the cost of considerably more intricate mathematical machinery (arenas, strategies, innocence conditions). In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly (via substitution), denotational semantics constructs it explicitly, and game semantics refines it via strategies. None of these approaches is first-order; [^16] Function spaces in denotational semantics, strategies in game semantics, and substitution in operational semantics. none eliminates the function space from the semantic domain; and none unifies denotational and operational semantics into a single first-order definition. All three also inherit $\beta$-reduction’s divergence on recursive programs: a naïve transitive closure on a cyclic graph, for instance, diverges in every framework.
@@
Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings. These models eliminate Scott topology but not the function space itself: they remain higher-order, encoding $D\to D]$ inside $D$ via injections, intersection types, or multiset relations. Graph models (Engeler, 1981; Plotkin, 1993) encode functions via a web injection on powerset lattices; Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$. Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the set of intersection types derivable for it; Dezani-Ciancaglini et al. (Dezani-Ciancaglini et al., 1998) showed that the BCD filter model’s $\lambda$-theory coincides exactly with Böhm tree equality. Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic via finite multisets; Ehrhard (Ehrhard, 2005) showed that a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category. All three families are semantic models: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure. None is a self-contained calculus with its own primitives into which the $\lambda$-calculus embeds. None is first-order, since the function space is encoded, not eliminated. And none provides a fully abstract denotational semantics of the lazy $\lambda$-calculus without higher-order machinery.
@@
Van Emden and Kowalski (van Emden and Kowalski, 1976) showed that the semantics of predicate logic as a programming language admits three equivalent characterizations: operational (SLD resolution), model-theoretic (least Herbrand model), and fixed-point (least fixed point of the immediate consequence operator $T_{P}$ on the powerset of ground atoms). Aczel (Aczel, 1977) gave a systematic treatment of inductive definitions via monotone operators on powerset lattices, providing the logical foundations that Datalog-style semantics rests on. Leroy and Grall (Leroy and Grall, 2006) treated divergence in big-step semantics via coinductive definitions; in inheritance-calculus, divergence already emerges as non-termination of recursive evaluation, so no separate coinductive formulation is needed. None of these prior frameworks has been applied to a first-order domain of paths to obtain a fully abstract semantics of the lazy $\lambda$-calculus without function spaces, substitution, or reduction rules. In particular, Van Emden and Kowalski’s $T_{P}$ computes $\mathrm{lfp}$ for recursive Datalog programs, but has not been used to give recursive $\lambda$-programs a fixpoint semantics in which they converge instead of diverging.
@@ 7.2. Inheritance, Objects, and Records
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records. Inheritance is composition of generators (functions from self to complete object) and wrappers (functions that modify generators), solved by a fixed-point construction. Cook’s key insight is that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods. This insight is one of the starting points of the present work. The model has four primitives (records, functions, generators, and wrappers) and requires a fixed-point combinator. Composition is asymmetric: a wrapper modifies a generator but not vice versa. The fixed-point construction relies on Scott continuity in a domain of functions. The resulting denotations are not directly executable, so a separate operational semantics is needed to define method dispatch and object construction at runtime.
@@
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses—functions from a superclass parameter to a subclass. Because mixin application is function composition, it is neither commutative nor idempotent; the C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order. Schärli et al. (Schärli et al., 2003) introduced traits, whose composition is commutative but rejects same-name conflicts (Ducasse et al., 2006). Both mechanisms operate at the method level: composing two definitions of the same name results in an override or a conflict, not a recursive merge of nested structure. Both presuppose the $\lambda$-calculus: method bodies are functions, and the composition mechanism itself is not Turing complete. Neither mixins nor traits are therefore deep-mergeable.
@@
Abadi and Cardelli (Abadi and Cardelli, 1996) treat objects as the primitive notion, but methods are functions that take self as a parameter and object extension is asymmetric. Boudol (Boudol, 2004) showed that self-referential records require an unsafe fixed-point operator whose well-definedness depends on the evaluation strategy. Harper and Pierce (Harper and Pierce, 1991) gave a record calculus with symmetric concatenation that rejects same-label composition; Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) introduced row polymorphism for records and variants in ML. All of these calculi operate on flat records: there is no recursive merging of nested structure, no lazy observation, and self-reference requires an explicit fixed-point combinator.

### 16 Mar 2026 v7

@@ 6.1. Emergent Phenomena
In developing real programs in MIXINv2, three phenomena arose from off-label use of the adopted design patterns. [^17] MIXINv2 programs adopt two standard patterns: data types follow a factory pattern isomorphic to Scala’s sealed trait with case classes; pattern matching uses Scott encoding. Both are named after Gang-of-Four terminology (Gamma et al., 1994) (factory, visitor, acceptance), with verbs replaced by nouns since the calculus has no methods. They can be seen in the source code (Yang, 2026) and the Case Study (Section 5). Below we explain why they arise from the three primitives.
@@
Single-valued computation and set-valued computation are the same mechanism. The semantic equations of Section 3 resolve every property to a set of paths, so operations defined on single values automatically distribute over all combinations of set-valued inputs (Section 5). When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise. The carrier is a trie rather than a flat relation, and observation is coinductive: the observer drives computation by querying progressively deeper paths. This contrasts with Datalog’s bottom-up materialization, but the relational structure of the computation is the same. Appendix F makes this correspondence precise with a systematic encoding of Datalog into inheritance-calculus, demonstrating that variable joins, constant matches, multi-body rules, and recursive rules are all expressible using only the three primitives of Section 2.
@@ 6.2. Semantic Alternatives
The preceding item established that inheritance-calculus is immune to nonextensibility. We adopt this as a design goal and ask whether an alternative design could share this property by surveying every candidate known to us.
@@
The analysis below assumes that the AST is a named tree:every edge from parent to child carries a label, so that every node is uniquely identified by the path from the root. For configuration languages this is naturally the case: every key in a JSON, YAML, or Nix attribute set is a label, and nesting produces paths. [^20] Nix is a functional language with first-class functions; its attribute sets alone do not form a named-tree DSL. The NixOS module system (NixOS Contributors, ), however, restricts the user-facing interface to nested attribute sets composed by recursive merging, and it is this DSL-level structure that naturally forms a named tree. For mainstream programming languages the AST is not directly a named tree, since anonymous expression positions and implicit scopes lack labels. After ANF transformation (Flanagan et al., 1993), however, every intermediate result is bound to a name, and the remaining anonymous positions [^21] For example, the body of a $\mathbf{let}$. can be assigned synthetic fresh labels such as $\mathrm{result}$ and $\mathrm{tailCall}$ in Section 4, yielding a named tree.
@@
Scala [^22] Scala is a multi-paradigm language. class, trait, object, method, block, conditional, match, while, throw, … variable, member access, extends, with, implicits, import, …
@@
Only the parameters placed by the author are extension points; the function body is closed. To make a new aspect extensible, the original function must be rewritten to accept an additional parameter. Any composition mechanism based on function application inherits this limitation, because $\beta$-reduction closes the function body. The Expression Problem (Wadler, 1998) is hard in such languages for the same reason: extensibility is opt-in. The frameworks that simulate immunity to nonextensibility (visitor pattern, object algebras (d. S. Oliveira and Cook, 2012), finally tagless interpreters (Carette et al., 2009)) are the machinery needed to work around this limitation.
@@
Removing or conditionally including definitions based on the presence of other definitions creates a dependency cycle: whether a definition exists depends on whether another definition exists, which in turn depends on the first. The recursive evaluation (Appendix A) cannot reach a fixed point when addition and removal alternate, causing divergence. This differs from the divergence inherent to Turing completeness: Turing-complete programs diverge because computation is unbounded, whereas conditional incorporation diverges because the immediate consequence operator $T_{P}$ oscillates rather than ascending monotonically.
@@
Single-target $\mathtt{this}$ is a promising candidate: the Single-Path Lemma (B.1) shows that for translated $\lambda$-terms, the frontier set $S$ in $\mathtt{this}$ (6) always contains exactly one path, so the $\lambda$-calculus embedding does not require multi-target resolution. However, single-target $\mathtt{this}$ is vulnerable to nonextensibility for a different reason. When multiple inheritance routes reach the same scope, single-target resolution must reject the situation or select one route. The NixOS module system (NixOS Contributors, ) raises a duplicate-declaration error (Appendix D). This rejects nested extensions where two independently authored modules define the same option, a composition that arises when solving the Expression Problem (Wadler, 1998) (Section 5).
Our first three implementations of MIXINv2 used standard techniques for $\mathtt{this}$ resolution: closure-based fixed points following Cook (Cook, 1989), and stack-based environment lookup using de Bruijn indices (de Bruijn, 1972). All three produced subtle bugs when inheritance introduced multiple routes to the same enclosing scope. The difficulty was structural. A closure captures a single environment, so a fixed-point combinator solves for a single $\mathtt{this}$ target. A stack-based environment is a linear chain where each scope level has one parent. Both data structures embed a single-valued assumption incompatible with the multi-target situation that multiple inheritance naturally produces (Appendix D). Equation (6) emerged from abandoning this assumption: it tracks a frontier set $S$ rather than a single current scope.
@@
Single-target $\mathtt{this}$ nested extensions rejected NixOS modules (NixOS Contributors, )
@@
We arrived at this semantics by trying each alternative in this section and finding that it failed: shallow merge could not compose nested extensions; priority merge introduced ordering dependencies; early binding froze references before composition could supply them; dynamic scope broke the substitution that the $\lambda$-calculus embedding relies on; single-target $\mathtt{this}$ rejected the multi-target compositions that the Expression Problem demands. Deep merge with multi-target late-binding $\mathtt{this}$ is what remained.
@@ 7.1. Computational Models and $\lambda$-Calculus Semantics
Schönfinkel (Schönfinkel, 1924) and Curry (Curry and Feys, 1958) showed that the $\lambda$-calculus reduces to three combinators $S$, $K$, II  [^24] The $I$ combinator is redundant since $I=SKK$. . The bracket abstraction that compiles $\lambda$-terms to SKI is mechanical but incurs exponential blow-up in term size; no new computational patterns emerge beyond those already present in the $\lambda$-calculus. Semantically, the two share the same models: Meyer (Meyer, 1982) showed that combinatory algebras satisfying five additional conditions are exactly lambda algebras. No independent semantic domain for SKI has ever been exhibited from which $\lambda$-calculus semantics could be derived. The direction has always been the reverse: Scott’s $D_{\infty}$ was built for the $\lambda$-calculus, and SKI was interpreted inside it afterwards. Dolan (Dolan, 2013) showed that the x86 mov instruction alone is Turing complete by exploiting addressing modes as implicit arithmetic on mutable memory; the resulting programs are correct but provide no abstraction mechanisms beyond those of the RAM Machine. The Turing Machine itself, while universal, lacks random access: reading a cell at distance $d$ requires $d$ sequential tape movements. Each of these models achieves Turing completeness within the semantic framework of an existing model, and no independent semantic domain has been exhibited for any of them.
@@
The semantics of the $\lambda$-calculus can be defined in several fundamentally different ways. Operational semantics takes $\beta$-reduction $(\lambda x.\,M)\,N\to M[N/x]$ as the primitive notion of computation (Church, 1936, 1941; Barendregt, 1984). This requires capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi (Abadi et al., 1991). It also requires a choice of evaluation strategy: call-by-name versus call-by-value (Plotkin, 1975) or the lazy strategy of Abramsky (Abramsky, 1990). Denotational semantics (Scott, 1970, 1976) interprets terms in a mathematical domain, but giving meaning to $\beta$-reduction requires a reflexive domain $D\cong[D\to D]$, which has no set-theoretic solution by Cantor’s theorem. Scott’s $D_{\infty}$ resolves this by constructing $D$ as an inverse limit of continuous lattices (Scott, 1972). The resulting semantics is adequate but not fully abstract (Plotkin, 1977; Milner, 1977): it identifies terms that observational equivalence distinguishes. Game semantics (Abramsky et al., 2000; Hyland and Ong, 2000; Nickau, 1994) achieves full abstraction for PCF and for the lazy $\lambda$-calculus (Abramsky and Ong, 1993; Abramsky and McCusker, 1995), requiring arenas, strategies, and innocence conditions. In each of these approaches, the function space $D\to D]$ is the central obstacle: operational semantics needs it implicitly via substitution. In contrast, denotational semantics constructs it explicitly, and game semantics refines it via strategies. Each of these approaches involves higher-order structure [^25] Function spaces in denotational semantics, strategies in game semantics, and substitution in operational semantics. and retains the function space in its semantic domain. The denotational and operational accounts remain separate formalisms in each case. All three inherit $\beta$-reduction’s divergence on recursive programs: a naïve transitive closure on a cyclic graph, for instance, diverges in every framework.
@@
Several families of $\lambda$-calculus models avoid the inverse-limit construction of Scott’s $D_{\infty}$ by working in set-theoretic rather than domain-theoretic settings. These models eliminate Scott topology but not the function space itself: they remain higher-order, encoding $D\to D]$ inside $D$ via injections, intersection types, or multiset relations. Graph models (Engeler, 1981; Plotkin, 1993) encode functions via a web injection on powerset lattices; Bucciarelli and Salibra (Bucciarelli and Salibra, 2008) proved that the greatest sensible graph theory equals the Böhm tree theory $\mathcal{B}$. Filter models (Barendregt et al., 1983; Coppo and Dezani-Ciancaglini, 1978) interpret a term as the set of intersection types derivable for it; Dezani-Ciancaglini et al. (Dezani-Ciancaglini et al., 1998) showed that the BCD filter model’s $\lambda$-theory coincides exactly with Böhm tree equality. Relational models (Ehrhard, 2005; Bucciarelli and Ehrhard, 2001) arise from the relational semantics of linear logic via finite multisets; Ehrhard (Ehrhard, 2005) showed that a standard reflexive object for the pure $\lambda$-calculus cannot exist in this category. All three families are semantic models: they interpret $\lambda$-terms in an external mathematical structure and validate $\beta$-reduction as an equation in that structure, rather than defining an independent calculus with its own primitives. Each encodes the function space rather than eliminating it, and each requires higher-order machinery to obtain full abstraction for the lazy $\lambda$-calculus.
@@
Van Emden and Kowalski (van Emden and Kowalski, 1976) showed that the semantics of predicate logic as a programming language admits three equivalent characterizations: operational [^26] Typically SLD resolution. , model-theoretic [^27] Based on least Herbrand models. , and fixed-point [^28] Namely, the least fixed point of the immediate consequence operator $T_{P}$ on the powerset of ground atoms. . Aczel (Aczel, 1977) gave a systematic treatment of inductive definitions via monotone operators on powerset lattices, providing the logical foundations that Datalog-style semantics rests on. Leroy and Grall (Leroy and Grall, 2006) treated divergence in big-step semantics via coinductive definitions, requiring a separate coinductive judgment to account for non-termination alongside the inductive evaluation judgment. Van Emden and Kowalski’s $T_{P}$ computes $\mathrm{lfp}$ for recursive Datalog programs over a first-order domain. Conventional $\lambda$-calculus semantics requires function spaces in the semantic domain, and recursive $\lambda$-terms ordinarily diverge under $\beta$-reduction; using $T_{P}$ over a first-order domain to give the $\lambda$-calculus a fixpoint semantics has not been explored.
@@ 7.2. Inheritance, Objects, and Records
Cook (Cook, 1989) gave the first denotational semantics of inheritance, modeling objects as recursive records. Inheritance is composition of generators  [^29] These are functions from self to complete object. and wrappers [^30] Wrappers are functions that modify generators. , solved by a fixed-point construction. Cook’s key insight is that inheritance is a general mechanism applicable to any form of recursive definition, not only object-oriented methods. This insight is one of the starting points of the present work. The model has four primitives (records, functions, generators, and wrappers) and requires a fixed-point combinator. Composition is asymmetric: a wrapper modifies a generator but not vice versa. The fixed-point construction relies on Scott continuity in a domain of functions. The resulting denotations are not directly executable, so a separate operational semantics is needed to define method dispatch and object construction at runtime.
@@
Bracha and Cook (Bracha and Cook, 1990) formalized mixins as abstract subclasses, that is, functions from a superclass parameter to a subclass. Because mixin application is function composition, it is neither commutative nor idempotent; the C3 linearization algorithm (Barrett et al., 1996) was later developed to impose a deterministic order. Schärli et al. (Schärli et al., 2003) introduced traits, whose composition is commutative but rejects same-name conflicts (Ducasse et al., 2006). Both mechanisms operate at the method level: composing two definitions of the same name results in an override or a conflict, not a recursive merge of nested structure. Both presuppose the $\lambda$-calculus: method bodies are functions, and the composition mechanism itself is not Turing complete. Neither mixins nor traits are deep-mergeable.
@@
Abadi and Cardelli (Abadi and Cardelli, 1996) treat objects as the primitive notion, but methods are functions that take self as a parameter and object extension is asymmetric. Boudol (Boudol, 2004) showed that self-referential records require an unsafe fixed-point operator whose well-definedness depends on the evaluation strategy. Harper and Pierce (Harper and Pierce, 1991) gave a record calculus with symmetric concatenation that rejects same-label composition; Cardelli (Cardelli, 1994) studied extensible records with subtyping; Rémy (Rémy, 1989) introduced row polymorphism for records and variants in ML. All of these calculi operate on flat records: there is no recursive merging of nested structure, no lazy observation, and $\mathtt{this}$ requires an explicit fixed-point combinator.